OpenStudy (anonymous):

solve for x A(3tx-2b)=c (dx-2)

OpenStudy (shamil98):

A(3tx-2b) = c(dx-2) A3tx - a2b = cdx - 2c a3tx = cdx + a2b -2c a3tx -cdx = a2b - 2c ............................................. I GIVE UP